Suppose that the New England Colonials baseball team is equally likely to win a game as not to win it. If 5 Colonials games are chosen at random, what is the probability that exactly 4 of those games are won by the Colonials?
Round your response to at least three decimal places.

Answer by stanbon(48568)

Suppose that the New England Colonials baseball team is equally likely to win a game as not to win it. If 5 Colonials games are chosen at random, what is the probability that exactly 4 of those games are won by the Colonials?
Round your response to at least three decimal places.
----------------------------------------------------------
P(win) = 1/2 = P(lose)
-------------------------
Your problem is binomial with n=5,p=1/2,x = 4
-----------------
Answer: P(x=4) = 5C4(1/2)^4(1/2) = 5(1/2)^5 = 5/32
====================================================
Cheers,
Stan H.